OftenIMITATED:

Determine whether the point (1, 5) is a solution to the system of equations. Explain your reasoning in complete sentences.

OftenIMITATED:

OftenIMITATED:

@shadow

ShoaibChaudhry:

look at the graph and look where it intersects And the intersecting line will be the answer

OftenIMITATED:

Can you possibly explain more thoroughly if possible @shoaibchaudhry ?

ShoaibChaudhry:

Im not sure on how exactly how 2 but that is how i was taught at school

ShoaibChaudhry:

But you can solve both equations

OftenIMITATED:

Oh okay,well thank you for trying ;anything and everything helps if you really were sincere about it.Have a good rest of your evening @shoaibchaudhry .

ShoaibChaudhry:

Np

OftenIMITATED:

@tonycoolkid21

J4ke:

solution for both fx and gx ?? or only gx ?

ShoaibChaudhry:

I think its both

ShoaibChaudhry:

1 is gx and 5 is fx

ShoaibChaudhry:

ans thet are asking is it write

snowflake0531:

There is only one intersection point point for g(x) and f(x), and (1,5) is not it

J4ke:

g(x)=3x+2 & our point (1,5); 1=x y=5 ; g(1)=3(1)+2 =5 so this point is solution for the gx equation

J4ke:

@snowflake0531 wrote:
There is only one intersection point point for g(x) and f(x), and (1,5) is not it
that's right but i think we should mention this note too

OftenIMITATED:

Thank you to everyone that helped and or is still helping,I really do appreciate it.Though,can anyone specifically clarify which is which to the question itself?

OftenIMITATED:

So would the answer to the question be to put down that the solution is

@j4ke wrote:
g(x)=3x+2 & our point (1,5); 1=x y=5 ; g(1)=3(1)+2 =5 so this point is solution for the gx equation

OftenIMITATED:

Or that (1,5) is not the point of intersection,I am probably overcomplicating the question ain't I..?

ShoaibChaudhry:

yes u are 1 is answer to gx and 5 is susposed to be answer to the other one

J4ke:

@oftenimitated wrote:
Or that (1,5) is not the point of intersection,I am probably overcomplicating the question ain't I..?
its not a solution for both equations its only for gx

ShoaibChaudhry:

but u think it is for boths since it sates the the point (1, 5) is a solution to the system of equations. this says for both

snowflake0531:

(1,5) is NOT a solution for both equations

snowflake0531:

the only solution is (0,2)

J4ke:

@shoaibchaudhry wrote:
but u think it is for boths since it sates the the point (1, 5) is a solution to the system of equations. this says for both
yeah I wish they screen shot the question too. that would be helpful

OftenIMITATED:

It was screenshotted,wasn't anyone able to notice...?

J4ke:

@oftenimitated wrote:
It was screenshotted,wasn't anyone able to notice...?
that was the graph not the question : )

OftenIMITATED:

surjithayer:

when we are to find the solution of a system of lines means we have to find a point which satisfy all the equations or it lies on all the lines. point of intersection lies on all the lines ,so it is the solution. Here (0,2) lies on both the lines at the same time . Hence it is the solution. (1,5) is not the solution.

J4ke:

So as @snowflake0531 said many times it's not a solution but notice that this point is a solution for g(x) your teacher expecting you to say something about it I don't think it's a coincidence its good to mention that too in your answer.

OftenIMITATED:

Oh okay so overall I would have to at least try my best and state in my own words of each value each person had helped input into the chat itself,you are saying @j4ke ?

snowflake0531:

do NOT ping me without a purpose please Jake, I appreciate the agreement but no

OftenIMITATED:

Anyways I highly appreciate everyone's help and sorry for overanalyzing,etc.Everyone,have a wonderful night.